OpenStudy (anonymous):

(2x^3+3x^2-19x-1) / (x-4)

OpenStudy (anonymous):

do you know how to use synthetic division?

OpenStudy (roadjester):

i was gonna ask that! *pouts*

OpenStudy (anonymous):

2 3 -19 -1 4 8 44 100 ________________________________ 2 11 25 99

OpenStudy (anonymous):

yes but its scketchy

OpenStudy (anonymous):

list the coefficients, which is the top line 2 3 -19 -1

OpenStudy (anonymous):

then since you are dividing by \[x-4\] put a 4 on the side

OpenStudy (anonymous):

bring down the 2 to the bottom line. 2 3 -19 -1 4 _____________________________ 2

OpenStudy (anonymous):

\[2\times 4 =8\] 2 3 -19 -1 4 8 ___________________________ 2

OpenStudy (anonymous):

\[8+3=11\] 2 3 -19 -1 4 8 ___________________________ 2 11

OpenStudy (anonymous):

\[4\times 11 = 44\] 2 3 -19 -1 4 8 44 ___________________________ 2 11

OpenStudy (anonymous):

\[-19+44=25\] 2 3 -19 -1 4 8 44 ___________________________ 2 11 25

OpenStudy (anonymous):

\[4\times 25=100\] 2 3 -19 -1 4 8 44 100 ___________________________ 2 11 25

OpenStudy (anonymous):

\[-1+100=99\] 2 3 -19 -1 4 8 44 100 ___________________________ 2 11 25 99

OpenStudy (anonymous):

you have divided a polynomial of degree 3 by a polynomial of degree 1 so your answer is a polynomial of degree 2 thus \[2x^2+11x+25+\frac{99}{x-4}\]